# DeDe: Detecting Backdoor Samples for SSL Encoders via Decoders

### v1

@@ 4.3 Methodology
$\vspace{-3mm}\begin{split}\mathcal{L}(h_{d},h_{e})&=\mathop{\mathbb{E}}_{\mathbf{x}\in\mathcal{D}}||\mathbf{x}-\hat{\mathbf{x}}||_{2}^{2},\\ {where}\quad\hat{\mathbf{x}}&=h_{d}\left(f_{\theta^{*}}(\mathbf{x}),h_{e}(M_{\alpha}\odot\mathbf{x})\right).\end{split}\vspace{-4mm}$ (1)
The testing stage of DeDe is calculating the loss the same as (1) on the test dataset. The only difference is that the choice of $\alpha$ in the testing stage is higher than in the training stage, in order to constrain the image space information. For example, $\alpha=0.9$ is used when training, then $\alpha=0.99$ is used in testing. It is to force the decoder more depending on the global embedding to detect malformed embedding with backdoor mapping. The testing metric is simple as follows:
@@
The reconstruction error $\ell\left(\mathbf{x}\right)$ is compared with a threshold $\tau$ to decide whether the inference is backdoor behavior, which is set to be 1.5×1.5$\times$1.5 × average training loss. The training algorithm of $h$ is summarized as Algorithm 1, where $\theta_{h}$ represents the trainable parameters in $h$.
@@ 5 Evaluation
We evaluate DeDe in two SSL paradigms: contrastive learning and CLIP. Their tasks are different as contrastive learning uses unlabeled image datasets for training encoder while CLIP uses image-text pair datasets to train a vision encoder and a text encoder. We use SimCLR [5] to perform contrastive learning and use the typical algorithm to execute CLIP. As we only consider attacks in CLIP’s vision encoder, the text encoder used is publicly available on OpenAI  [^1] https://github.com/openai/CLIP/.
@@ 5.1 Experiment Set-up
Table 1: Upstream Detection Performance
@@
Attack models. We choose several representative and SOTA attacks in SSL to acquire the victim encoder. They are BadEncoder [22], CTRL [24], DRUPE [35], CLIP-Backdoor [4] and BadCLIP  [2]. The first three are attacks on CL and the last two are attacks on CLIP. BadEncoder, DRUPE, and CLIP-Backdoor use patch-like triggers while CTRL and BadCLIP use invisible(imperceptible) triggers. We use their public repository to recover the attacks and evaluate their functionality on a uniform downstream baseline to ensure the clean performance and the backdoor performance, except for CLIP Backdoor for unavailability, which we use the implementation from [12].
@@
Detection models. There are two SOTA detection methods in SSL to our knowledge: DECREE[13] and ASSET[29]. For settings, ASSET’s training requires a small but clean dataset, referred to as the base dataset, while their test criteria are identical to ours, which is convenient to compare with. DECREE optimizes the trigger pattern with a shadow dataset and a given model, subsequently employing a criterion on the optimized trigger pattern to identify backdoor behaviors. However, it fails to detect any backdoor samples during testing. Therefore, we compare the optimized trigger and the poisoned dataset, and calculate their similarity to detect any backdoor samples. Specifically, we assess the $l_{2}$-norm between the optimized trigger and the image data located within the trigger region, and use the norm as the error measurement. Utilizing this error for each data point enables us to compute all scores accordingly. For DECREE, we put * on detection scores to denote a successful detection of the model backdoor.
@@
Datasets. As different stages of training and testing are involved, we divide the dataset into the following categories: pre-train dataset, shadow dataset, memory dataset, DeDe training dataset, downstream training dataset and downstream testing dataset.
@@
Since we mainly use pre-trained model checkpoints for our implementation, pre-train dataset is omitted. For shadow dataset, we follow attacks’ public implementations, so we simply mention their datasets usage here. CIFAR-10 [23] is modified as the backdoor training dataset for BadEncoder, DRUPE and CTRL and ImageNet [9] is used for backdoor injection in CLIP Backdoor and BadCLIP, whose clean models are trained on the CC3M dataset [34]. For memory dataset, it is the clean version of shadow dataset. In ASSET, it uses a subset of memory dataset. For DeDe training dataset, we either use a subset from the same dataset in backdoor training, i.e. CIFAR-10 for CL and ImageNet for CLIPs, or a out-of-distribution dataset STL-10 [8] consistently. As we mentioned in Section 3.1, whether the downstream training data is poisoned is crucial to some attacks. We consider datasets CIFAR-10, GTSRB [20] and SVHN [28], and two cases of each dataset by clean and poisoning with $1\%$ backdoor samples. The downstream testing datasets are simply the clean test dataset and backdoor test dataset for evaluation.
@@
No Def. Poison 90.5 18.8 86.3 100.0 73.9 95.4 85.6 100.0 79.4 100.0 82.8 98.3
No Def. Clean 90.7 10.2 86.4 99.9 74.0 63.8 85.8 96.3 79.7 98.4 83.1 11.7
@@
Metrics. To compare the detection performance, we mainly use three related metrics: true positive rate (TPR), false positive rate (FPR) and area under the ROC curve (AUC score). The detection of backdoor samples is treated as a binary classification problem. Typically, the poisoned samples are in the minority to evade detection and scrutiny. TPR, also referred to as Recall, is calculated by $TP/(TP+FN)$, which measures the accuracy of correctly detected backdoor samples. FPR compensates the TPR by reflecting the clean samples that are misclassified as backdoor, formulated by $FP/(FP+TN)$. For imbalanced binary data distribution, it is common to measure TPR, FPR as used in [13, 29]. But the (TPR,FPR) pair only reflect a single point on the ROC curve with a fixed threshold. To authentically represent the detection result independently from the threshold, the ROC curve is plotted by sliding the threshold. The AUC score measures the area under ROC curve and distributes in $(0,1]$, where $1$ stands for a perfect classifier. Since TPR and FPR represent a trade-off score depending on a threshold, we present the default result calculated by the scikit-learn library [30].
@@
Following the prior work [29], we evaluate the detection performance also by the downstream tasks after removing the detected points. The metrics are the clean accuracy (CA) for the main task and the attack success rate (ASR) for the backdoor task. CA reflects the model utility of embedding, which means the higher the CA is, the better the encoder performs. ASR reflects the backdoor effectiveness, which means the higher the ASR is, the more successful the backdoor is. As explained in the dataset part, the downstream setting in attacks are relevant. So for detection settings, we set all downstream training dataset to be poisoned by $1\%$ as the challenge for detectors. Detection methods are used to filter the training and testing dataset before actual training and testing. the detected testing data are counted as negative samples in evaluating ASR. In this case, high CA indicates enough clean samples are kept for the main task training and low ASR indicates backdoor samples are successfully filtered by the algorithm.
@@
Model architectures. As DeDe has no dependency on the victim encoder, we use default backbone architectures for encoders, namely ResNet-18 for BadEncoder, DRUPE, CTRL, and ViT-B/16 for CLIPs. For DeDe, we use the lightweight MAE decoder developed from ViT-B/16 [1] as described in  [19]. Our default setting only has 6 blocks for the encoder and 4 blocks for the decoder, which has a width of 512 for ResNet and 1024 for ViT. The default training mask ratio is set to be $0.9$. The default patch size is set to be 4 for CL with image size $32\times 32$ and CLIP with image size $224\times 224$.
@@
Auxiliary Explanation. Recall in Section 3.1, we define a good backdoor in SSL setting to be imitative mapping behavior towards the target embeddings, which is independent of the downstream classifier. CTRL falls short in such a definition as its end-to-end backdoor training. For this reason, it embeds the triggered inputs in the out-of-distribution region of the embedding space, as illustrated in Figure 4(a). One direct consequence is that, if we use a clean training dataset in the downstream task, its ASR drops in the training iterations, which we show in Figure 4(b). Despite this, we report the ASR in the first epoch in Table 2 for CTRL. When training DeDe for CTRL, we address the issue of out-of-distribution embeddings by utilizing random noise images to fill the out-of-distribution regions, where we define the region by a norm ball with a radius of 0.99 quantiles. Furthermore, the training masking ratio for CTRL is set at $0.75$ to provide additional support from the image space, as the embeddings are significantly sparse.
@@
Figure 4: CTRL Attack
@@ 5.2 Experiment Results
We assess detection performance in two scenarios: one with a balanced data distribution in which 50% of the data is triggered, and another with an unbalanced data distribution where only 1% of the data is triggered. For ASSET, a clean subset is generated for their model training, whose size is $5$k for CL and $500$ for CLIP. We observe that ASSET’s performance is unstable when using nested offsets (epoch number is set to $10$), the best epoch is reported. See the supplementary material for omitted results.
@@
In Table 2, as we discussed in Section 3.1, even the clean baseline shows an increase in ASR when the downstream task is trained on a contaminated dataset. Consistently, the no-defense baselines always have higher attack success rates when the poisoned training dataset is used instead of a clean one. This trend is particularly evident in the ASR rate for CTRL, highlighting its dependence on end-to-end testing. Regarding ASSET, it demonstrates defensive utility against BadEncoder, CTRL, and DRUPE; however, when compared to our approach, its defense capability is approximately 40% lower in terms of attack success probability. Furthermore, ASSET completely fails to defend against CLIP attacks, whereas we achieve a 100% success rate in defending against CLIP Backdoor and a 70% success rate in defending against BadCLIP. Though ASSET demonstrates the detection ability to detect BadCLIP to some extent in Table 1, but CLIP backdoor is effective as long as a few images pass the sifting.
@@
We also present two representatives of DRUPE and CLIP-backdoor for the decoding results in Figure 5. In the images for DRUPE, it is noticeable that all reconstructed images have the trigger pattern, which suggests their embeddings are highly corrupted towards the backdoor distribution leading to high ASR presented in Table 1. Also as a result, our reconstruction effectively reverses the corrupted embeddings, leading to a significant error when compared to the input images. For CLIP images, the reversed images lose grid-shaped pixel information due to the corrupted embeddings. Although the remaining pixels of the reconstruction are similar, it still results in a sufficiently large loss that distinguishes backdoor samples in the test dataset.
@@ 6 Conclusion
In this paper, we propose an SSL backdoor detection method DeDe that has no reliance on knowledge of the training dataset or the trigger type. Our empirical study demonstrates the DeDe’s detection capability among different learning paradigms of CL and CLIP, as well as more stealthy backdoor attacks, hence, improving the scarcity and limitations in SSL backdoor countermeasures.

### v2

@@ 4.3 Methodology
$\vspace{-2mm}\begin{split}\mathcal{L}(h_{d},h_{e})&=\mathop{\mathbb{E}}_{\mathbf{x}\in\mathcal{D}}||\mathbf{x}-\hat{\mathbf{x}}||_{2}^{2},\\ {where}\quad\hat{\mathbf{x}}&=h_{d}\left(f_{\theta^{*}}(\mathbf{x}),h_{e}(M_{\alpha}\odot\mathbf{x})\right).\end{split}$ (1)
The testing stage of DeDe is calculating the loss the same as (1) on the test dataset. The only difference is that the choice of $\alpha$ in the testing stage is higher than in the training stage, in order to constrain the image space information. For example, $\alpha=0.9$ is used when training, then $\alpha=0.99$ is used in testing. It is to force the reconstruction more depending on the global embedding to detect malformed ones. The testing metric is simple as follows:
@@
The reconstruction error $\ell\left(\mathbf{x}\right)$ is compared with a threshold $\tau$ to decide whether the inference is backdoor behavior. To tolerate normal reconstruction error, the threshold is set to be 1.5×1.5$\times$1.5 × average training loss. The training algorithm is summarized as Algorithm 1, where $\theta_{h}$ represents the trainable parameters in $h_{d},h_{e}$.
@@ 5 Evaluation
We evaluate DeDe in representative SSL paradigms: contrastive learning and CLIP. Their tasks are different as contrastive learning uses unlabeled image datasets for training encoder while CLIP uses image-text pair datasets to train a vision encoder and a text encoder. We use SimCLR [5] to perform contrastive learning and use common training for CLIP [32]. As we only consider attacks in the vision encoder, the used text encoder is publicly available on OpenAI  [^1] https://github.com/openai/CLIP/. Our code is publicly available at this repository https://github.com/tardisblue9/DeDe.
@@ 5.1 Experiment Set-up
Table 1: Upstream Detection Performance (* is used to denote a successful detection for DECREE.)
@@
Attack models. We choose several representative and SOTA attacks in SSL to acquire the victim encoder. They are BadEncoder [23], CTRL [25], DRUPE [36], CLIP-Backdoor [4] and BadCLIP  [27]. The first three are attacks on CL and the last two are attacks on CLIP. BadEncoder, DRUPE, and CLIP-Backdoor use patch-like triggers while CTRL and BadCLIP use invisible(imperceptible) triggers, specifically frequency-based and adaptive triggers respectively. We use their public repositories to recover the attacks and evaluate their functionality on a uniform downstream baseline to ensure the clean performance and the backdoor performance, except for CLIP Backdoor for unavailability, which we use the implementation from [12].
@@
Detection models. There are two SOTA detection methods in SSL to our knowledge: DECREE[13] and ASSET[30]. For settings, ASSET’s training requires a small but clean dataset, referred to as the base dataset, while their test criteria are identical to ours, which is convenient to compare with. DECREE optimizes the trigger pattern with a shadow dataset, subsequently employing a criterion on the optimized trigger pattern to identify backdoor behaviors. However, it fails to detect any backdoor samples during testing. Therefore, we compare the optimized trigger and the poisoned dataset, and calculate their similarity to detect any backdoor samples. Specifically, we assess the $\ell_{2}$-norm between the optimized trigger and the image data located within the trigger region, and use the norm as the error measurement. Utilizing this error for each data point enables us to compute all scores accordingly. For DECREE, we use * on scores for successful detection of the model backdoor.
@@
Datasets. As different stages in the pipeline are involved, we divide the dataset into the following categories: pre-train dataset, shadow dataset, memory dataset, DeDe training dataset, downstream training/testing dataset. Unified augmentation is used to rectify different color distributions.
@@
Since we mainly use pre-trained model checkpoints for our implementation, pre-train dataset is omitted. For shadow dataset, we follow attacks’ public implementations, so we simply mention their datasets usage here. CIFAR-10 [24] is modified as the backdoor training dataset for BadEncoder, DRUPE and CTRL and ImageNet [9] is used for backdoor injection in CLIP Backdoor and BadCLIP, whose clean models are trained on the CC3M dataset [35]. For memory dataset, it is the clean version of shadow dataset. In ASSET, it uses a subset of memory dataset. For DeDe training dataset, we either use a subset from the same dataset in backdoor training, i.e. CIFAR-10 for CL and ImageNet for CLIPs, or an out-of-distribution dataset (OoD) STL-10 [8] consistently. As we mentioned in Section 3.1, whether the downstream training data is poisoned is crucial to some attacks. We consider datasets CIFAR-10, GTSRB [20] and SVHN [29], and two cases of each dataset by clean and poisoning with $1\%$ backdoor samples. The downstream testing datasets are simply the clean test dataset and backdoor test dataset for evaluation.
@@
No Def. Poison dataset 90.5 18.8 86.3 100.0 73.9 95.4 85.6 100.0 79.4 100.0 82.8 98.3
No Def. Clean dataset 90.7 10.2 86.4 99.9 74.0 63.8 85.8 96.3 79.7 98.4 83.1 11.7
@@
Metrics. To compare the detection performance, we mainly use three related metrics: true positive rate (TPR), false positive rate (FPR) and area under the ROC curve (AUC score). The detection of backdoor samples is treated as a binary classification problem. Typically, the poisoned samples are in the minority to evade detection and scrutiny. TPR, also referred to as Recall, is calculated by $TP/(TP+FN)$, which measures the accuracy of correctly detected backdoor samples. FPR compensates the TPR by reflecting the clean samples that are misclassified as backdoor, formulated by $FP/(FP+TN)$. For imbalanced binary data distribution, it is common to measure TPR, FPR as used in [13, 30]. But the (TPR,FPR) pair only reflects a single point on the ROC curve with a fixed threshold. To authentically represent the detection result independently from the threshold, the ROC curve is plotted by sliding the threshold. The AUC score measures the area under ROC curve and distributes in $(0,1]$, where $1$ stands for a perfect classifier. Since TPR and FPR represent a trade-off score depending on a threshold, we present the default result by the scikit-learn library [31].
@@
Following the prior work [30], we evaluate the detection performance also by the downstream tasks after removing the detected points. The metrics are the clean accuracy (CA) for the main task and the attack success rate (ASR) for the backdoor task. CA reflects the model utility of embedding, which means the higher the CA is, the better the encoder performs. ASR reflects the backdoor effectiveness, which means the higher the ASR is, the more successful the backdoor is. As explained in the dataset part, the downstream setting in attacks is relevant. So for detection settings, we set all downstream training dataset to be poisoned by $1\%$ as the challenge for detectors. Detection methods are used to filter the training and testing dataset before actual training and testing. The detected testing data are counted as negative samples in evaluating ASR. In this case, high CA indicates enough clean samples are kept for the main task training and low ASR indicates backdoor samples are successfully filtered by the algorithm.
@@
Model architectures. As DeDe has no dependency on the victim encoder, we use default backbone architectures for encoders, namely ResNet-18 for BadEncoder, DRUPE, CTRL, and ViT-B/16 for CLIPs. For DeDe, we use the lightweight MAE decoder developed from ViT-B/16 [1] as described in  [19]. Our default setting only has 6 blocks for the encoder and 4 blocks for the decoder, which has a width of 512 for ResNet and 1024 for ViT. The default training mask ratio $\alpha$ is set to be $0.9$ with an exception for CTRL. The default patch size is chosen between 4 to 32 for image size from $32\times 32$ in CL and $224\times 224$ in CLIP. For datasets inconsistent with the image input size (such as STL-10), up/down sampling are used respectively.
@@
Auxiliary Explanation. Recall in Section 3.1, we define a good backdoor in SSL to be imitative mapping behavior towards the target embeddings, which is independent of the downstream classifier. CTRL falls short in such a definition for its end-to-end backdoor training. Hence, it embeds the triggered inputs in the out-of-distribution region across the embedding space as illustrated in Figure 4(a). One direct consequence is that, if we use a clean training dataset in the downstream task, its ASR drops in the training iterations as in Figure 4(b). Despite this, we report the highest ASR for CTRL in Table 2. When training DeDe against CTRL, we address the issue by filling the out-of-distribution regions with random noise image embeddings, where we define the region outside a norm ball with radius of 0.99 quantiles in embedding norms. Furthermore, the training masking ratio for CTRL is set at $0.75$ to provide additional support from the image space, as the embeddings are significantly sparse.
@@
Figure 4: Discussion for CTRL Attack
@@ 5.2 Experiment Results
We assess detection performance in two scenarios: one with a balanced data distribution in which 50% of the data is triggered, and another with an unbalanced data distribution where only 1% of the data is triggered (see Appendix). For ASSET, a clean subset is sampled for their model training, whose size is $5$k for CL and $500$ for CLIP. We observe that ASSET’s performance is unstable when using nested offsets (epoch number is set to $10$), the best epoch is reported.
@@
In Table 2, as we discussed in Section 3.1, even the clean baseline shows an increase in ASR when the downstream task is trained on a contaminated dataset. Consistently, the no-defense baselines (No Def.) always have higher attack success rates when the poisoned training dataset is used instead of a clean one. This trend is particularly evident in the ASR rate for CTRL and BadCLIP, highlighting their dependence on end-to-end testing. Regarding ASSET, it demonstrates defensive utility against BadEncoder, CTRL, and DRUPE; however, when compared to our approach, its defense capability is approximately 40% lower in terms of attack success probability. Furthermore, ASSET completely fails to defend against CLIP attacks, whereas we achieve a 100% success rate in defending against CLIP Backdoor and a 70% success rate in defending against BadCLIP. Though ASSET demonstrates the detection ability to detect BadCLIP to some extent in Table 1, but CLIP backdoor is effective as long as a few images pass the sifting.
@@
We also present two representatives of DRUPE and CLIP-backdoor for the decoding results in Figure 5. In the images for DRUPE, it is noticeable that all reconstructed images have the trigger pattern, suggesting their embeddings are highly corrupted towards a backdoor distribution leading to high ASR presented in Table 1. Also as a result, our reconstruction effectively reverses the corrupted embeddings, leading to a significant error when compared to the input images. For CLIP images, the reversed images lose grid-shaped pixel information due to the corrupted embeddings. Although the remaining pixels of the reconstruction are similar, it still results in a sufficiently large loss that distinguishes backdoor samples in the test dataset. More results for different dataset, unbalanced data distribution and reconstruction demonstrations for frequency-based/imperceptible attacks are presented in Appendix.
@@ 6 Conclusion
In this paper, we propose an SSL backdoor detection method DeDe that has no reliance on knowledge of the training dataset or the trigger type and is non-invasive to the victim encoder. Our empirical study demonstrates the DeDe’s detection capability among different learning paradigms of CL and CLIP, as well as more stealthy backdoor attacks, hence, improving the scarcity and limitations in SSL backdoor countermeasures.
Acknowledgement
This work is in part supported by the Fundamental Research Funds for the Central Universities (Grant No. 2242024k30059).
Appendix
Appendix A Additional Visualizations
There is a common property in embedding distribution: images with similar features tend to cluster together in SSL. We present t-SNE plots for better visualization of different SSL paradigms in Fig.6, including CL, CLIP, MAE, and auto-regressive model. With no exception, samples from the same classes are clustered in the embedding space. It serves a fundamental fact in various SSL models, which we have leveraged when designing DeDe.
Figure 6: t-SNE visualization of embeddings in different SSLs.
Appendix B Additional Experimental Results
  • Table 3: Upstream detection results for unbalanced data.
  • Table 4: Downstream defense performance for GTSRB.
  • Table 5: Downstream defense performance for SVHN.
  • Table 6: The reconstruction examples and error distributions, in which different parameter settings are presented.
As a reminder, the balanced dataset presented in the paper is half poisoned (50%) test dataset of size $10000$. Table 3 presents the result of a slightly poisoned (1%) test dataset of the same size. ASSET’s training dataset is kept with the same poisoning rate as the test dataset. In the context of upstream detection, the observed trend aligns with what is presented in balanced data. Although DECREE demonstrates low AUC scores, it successfully identifies backdoor attacks in BadEncoder and CLIP-Backdoor. ASSET is capable of detecting BadEncoder and CTRL attacks but is ineffective against the stealthy DRUPE and CLIP-Backdoor attacks. Notably, ASSET shows strong performance in BadCLIP, suggesting its sensitivity to the selected poisoning rate. In contrast, DeDe maintains consistent performance across all attacks. The downstream defense performance shows a decrease relative to the performance on CIFAR10, as reported in the main text. In comparison, ASSET demonstrates defense capabilities against CL attacks but is ineffective against CLIP attacks. Conversely, DeDe achieves an approximately $40\%$ improvement over all CL attacks and consistently defends against CLIP attacks. Although the defense performance in BadCLIP is not as strong as in other cases, it still reduces the attack success rate to 30% in both scenarios.
In Table 6, we present reconstruction examples for all attacks using our method, DeDe. We use the same samples to present consistent visualization, so the images for CLIP and BadCLIP are up-sampled to $224\times 224$ for demonstration. To present the robustness in comparison to ASSET, we present the error histograms of both methods, as they are both unsupervised techniques for detecting backdoor samples by computing losses. ASSET effectively distinguishes between clean and backdoor samples in the case of BadEncoder, successfully separating the two modes. However, its performance declines when it struggles to differentiate the backdoor samples in the first place, resulting in a mixing of the two modes. In contrast, while DeDe does not push the discerned samples further, it demonstrates sufficient strength to create two distinct modes, allowing for the use of a threshold to filter samples. It is also worth noting, that the results of DeDe are generally stable for different choices of patch size and masking ratio. It is reasonable to choose masking ratio in the range of $[0.75,0.95]$, which is supported by our testing results.
DeDe training overhead. In training the DeDe decoder, the given encoder(poisoned) is frozen for inference. Learning is on ViT-B/16 model as the decoder. We use a machine with Intel(R) Xeon(R) Gold 5118 CPU@2.30GHz and NVIDIA GeForce RTX 4090 GPU. Taking DRUPE as an example, max epoch is set to 200 and DeDe training dataset size is 50k. The total run time is 1hr’11m’30s, which is around 20 s/epoch. The training time is generally consistent in different attacks while experiments with $224\times 224$ image dataset take a bit longer.
Table 3: Upstream Detection Performance for unbalanced data.
BadEncoder CTRL DRUPE CLIP Backdoor BadCLIP
DECREE 66.0 42.1 0.336 - - - 68.0 49.8 0.366 70.0 49.7 0.363 - - -
ASSET 100.0 25.0 0.901 31.1 91.0 0.799 94.8 27.6 0.858 54.4 47.8 0.555 100.0 25.0 0.943
DEDE 92.0 8.3 0.978 90.0 19.3 0.898 98.5 3.1 0.998 100.0 0.0 1.0 84.0 19.4 0.903
DEDE OOD 96.5 8.2 0.983 81.0 19.3 0.853 95.5 8.3 0.979 100.0 0.0 1.0 88.0 19.3 0.936
Table 4: Downstream Performance for GTSRB.
No Attack BadEncoder CTRL DRUPE CLIP Backdoor BadCLIP
No Def. Poison 85.14 19.16 82.03 99.36 73.9 91.57 85.6 99.46 73.76 97.24 76.99 98.59
No Def. Clean 85.25 10.85 82.26 97.27 67.48 64.26 81.60 97.03 74.51 98.05 79.74 11.69
ASSET - - 83.24 54.67 67.45 50.51 80.85 54.07 73.92 96.42 76.19 98.38
DEDE - - 82.86 2.99 68.61 4.43 80.46 0.91 74.41 2.21 76.20 30.60
Table 5: Downstream Performance for SVHN.
No Attack BadEncoder CTRL DRUPE CLIP Backdoor BadCLIP
No Def. Poison 84.22 17.43 79.97 99.77 73.9 88.22 85.6 97.41 73.03 98.44 75.31 98.01
No Def. Clean 85.38 10.34 78.25 99.70 67.84 63.29 77.56 96.86 73.89 97.23 75.45 11.57
ASSET - - 81.09 56.86 66.52 48.95 79.26 53.45 70.50 99.39 73.88 95.80
DEDE - - 79.28 1.53 64.70 4.78 80.03 1.85 73.93 1.82 75.12 30.51
[…]
BadEncoder CTRL DRUPE CLIP BadCLIP
Reconstruction Examples
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
DeDe Dist.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Asset Dist.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]